OpenStudy (anonymous):

Find the area of the region that is enclosed between (y = 6x^{2}-x^{3}+x) and (y =x^{2}+5x)

OpenStudy (amistre64):

do we know where they intersect yet?

OpenStudy (anonymous):

I see that they intersect at 0 and (1,3)

OpenStudy (anonymous):

Sorry, more like (1,5)

OpenStudy (amistre64):

there are 3 points of intersection tha ti see

OpenStudy (amistre64):

6x^2-x^3+x =x^2+5x; x=0 is obvious so -x^2+6x+1 =x+5 0 = x^2 -5x+4 0 = (x-4)(x-1) , x=1 and x=4 are the other intersections

OpenStudy (amistre64):

so we need to split out integral into 2 intevals and add the results 0 to 1 and 1 to 4

OpenStudy (amistre64):

you have any thoughts on what to integrate?

OpenStudy (anonymous):

I'm guessing one integral will have upper and lower limits of those two intersections, whilst integrating the equations they ask of us?

OpenStudy (amistre64):

seems fair lets clarify a bit

OpenStudy (amistre64):